A chemist needs 130 ml of a 31% acid solution, but has only 7% and 46% solutions available. How many ml of the 46% solution should be added to get the desired solution?
Answer by richwmiller(17219)

We need a total of 130 mL at 31%
We want to know how many mL at 7% to mix with 46%
a+b=130,
0.07*a+0.46*b=0.31*130
a=130-b
0.07*(130-b)+0.46*b=40.3
9.1-0.07b+0.46*b=40.3
0.39*b=31.2
b=80
a=130-b
a=50 mL at 7%
b=80 mL at 46%
check
0.07*50+0.46*80=0.31*130
3.5+36.8=40.3
40.3=40.3
ok
codemx1